# Concept-Aware LoRA for Domain-Aligned Segmentation Dataset Generation

### v1

@@ Abstract
Work done during an internship at Qualcomm AI Research. Corresponding author.   Qualcomm AI Research is an initiative of Qualcomm Technologies, Inc.
@@ 1 Introduction
The amount of labeled data is crucial for achieving high performance in semantic segmentation. However, acquiring diverse image samples, especially in rare or complex scenarios, and providing pixel-wise annotations are labor-intensive and time-consuming. Recent advances in text-to-image (T2I) generation models [41, 43, 37] have significantly improved image quality, enabling their use for data creation in perception tasks such as segmentation with minimal human effort. Recent studies [59, 2, 51, 52] leverage these models, such as Stable Diffusion [41], pretrained on large-scale datasets [46]. By utilizing the rich generative features from T2I models, label generation modules can be trained with minimal labeled data to parse semantic regions. Furthermore, these models also provide controllability through text input, allowing for the generation of underrepresented distributions. These approaches have proven particularly effective in augmenting labeled datasets [59, 16, 1, 51], addressing data scarcity, and creating desired distributions, such as long-tailed class distributions [47] or adverse weather condition [23].
There are two primary challenges in generating segmentation datasets using image generation models: 1) aligning with the target domain, and 2) generating informative datasets beyond existing training data. However, previous approaches have overlooked key aspects of these issues. Existing methods [59, 29, 2, 35] typically produce domain-aligned but non-informative samples, as they train image generation models from scratch using only segmentation datasets, resulting in images that closely resemble the training data. To overcome this, leveraging T2I models such as Stable Diffusion, pretrained on large-scale datasets, is crucial for generating more diverse and informative samples. However, recent methods [51, 52, 56, 23] often use these pretrained models without fine-tuning for segmentation tasks, leading to poor alignment with the target domain.
Full fine-tuning or Low-Rank Adaptation (LoRA) [21] of pretrained T2I models are potential solutions to the challenges discussed above. However, as shown in Fig. 1, even LoRA fine-tuning often overfits and memorizes training data, limiting the generation of diverse and informative samples. This occurs because the model learns all concepts (e.g., viewpoint, style, object shape, and layout) existing in the training data. Therefore, we need a method to selectively learn only the necessary concepts (e.g., viewpoint or style) for domain alignment, while leveraging the pretrained knowledge of T2I models to generate informative samples.
The problem setting (e.g., in-domain or domain generalization (DG)) determines the target domain, affecting the concepts the model should learn during fine-tuning. In in-domain settings, the model needs to learn the style of the training data. However, in DG settings, where the target domain is unseen, learning only the viewpoint is more effective. For example, if the target domain is ACDC [45], which contains driving scenes under adverse weather conditions, learning all concepts from Cityscapes [9], containing only clear-day driving scenes, is suboptimal. Instead, the model should generate driving scenes (domain-aligned) under diverse weather conditions (informative) by learning only viewpoints from Cityscapes, excluding its clear-day style. However, as shown in Fig. 1, a pretrained model often fails to generate proper driving viewpoints, while a LoRA fine-tuned model generates only the clear-day style, even when ‘foggy’ or ‘night-time’ conditions are specified in the prompts. Therefore, a method that selectively learns only viewpoints from Cityscapes is necessary.
To the best of our knowledge, this is the first work to address these issues comprehensively. We propose Concept-Aware LoRA (CA-LoRA) that updates only the weights related to desired concepts (e.g., viewpoint or style) while preserving the rest to leverage the knowledge of pretrained T2I models. Our approach enables the model to learn only the key concepts needed for aligning with the target domain, resulting in generated images that are not only well-aligned but also more diverse and informative. For example, if the desired concept is driving-scene viewpoints, the model learns that viewpoint alone and generates images that extend beyond the original training data by incorporating various styles, object shapes, and layouts. Additionally, the text controllability of T2I models enables the generation of specific styles from user input, making them highly effective in DG setting [7, 20, 19], which require diverse conditions such as adverse weather or varying illumination.
@@
We demonstrate the effectiveness of our approach for urban-scene segmentation by comparing it with state-of-the-art dataset generation methods [51, 23, 5, 4] in both in-domain and DG settings. Our contributions are threefold:
@@
    We propose Concept-Aware LoRA, a novel fine-tuning method that selectively identifies and updates only the weights related to the necessary concepts (e.g., viewpoint or style) for domain alignment, reducing overfitting and preserving pretrained knowledge.
@@
    Our method achieves state-of-the-art performance across various tasks, with improvements of +2.30 mIoU in few-shot, +1.34 mIoU in fully supervised settings on Cityscapes, and +1.53 mIoU in DG benchmarks (ACDC, Dark Zurich, BDD100K, Mapillary Vistas), along with strong results on Pascal-VOC. It consistently generates higher-quality image-label pairs than existing methods.
Figure 2: Overview of the proposed framework for generating an urban-scene segmentation dataset by learning the Cityscapes viewpoint. The process consists of four stages: (1) identifying sensitive weights for a specific concept, (2) selectively fine-tuning them with LoRA, (3) training a label generator using features from T2I model, and (4) generating diverse image-label pairs with augmented prompts.
@@ 2.2 Segmentation Dataset Generation
Recent studies have focused on leveraging the pretrained knowledge of T2I models [51, 52, 34, 4], but often overlook the alignment of the generated images with the target domain (e.g., style, viewpoint). This paper explores the impact of fine-tuning T2I models for segmentation dataset generation, focusing on ensuring better image alignment.
Figure 3: Overview of measuring concept sensitivity. (a) We design the concept loss ($\mathcal{L}_{\text{Concept}}$) with the concept-augmented captions ($c_{\text{Aug}}$), and the original diffusion loss ($\mathcal{L}_{\text{Diffusion}}$) with the added noise $\epsilon$. The concept-augmented captions can be changed according to the desired concept (e.g., style, viewpoint). (b) While each concept gradient represents the reaction of the concept, it has to be normalized with the original diffusion gradient to assess the increased ratio of each layer.
@@ 3.1 Overall Framework
Fig. 2 shows our dataset generation framework. First, we identify weights sensitive to the desired concept to be learned from the source dataset (Stage 1). Next, Concept-Aware LoRA selectively fine-tunes the top-$k$sensitive weights of the T2I model (Stage 2). Then, a label generator is trained on a labeled dataset using rich image generative features from the T2I model (Stage 3). Finally, diverse image-label pairs are generated with augmented prompts according to the problem settings (Stage 4).
@@ 3.2 Identify Sensitive Weights to Desired Concept
To identify the sensitive weights for a specific concept (e.g., style, viewpoint), we design an objective function, Concept loss ($\mathcal{L}_{\text{Concept}}$), which can be flexibly changed according to the desired concept. As shown in Fig. 3 (a), Concept loss is applied to the noisy image to enforce modifying the concept. For example, when the Concept loss forces the T2I model to modify style (e.g., photorealistic $\rightarrow$ sketch), its gradient can be used to identify style-sensitive weights.
First, for the Concept loss input, we generate a few images $x_{0}$ using the T2I model $\Phi_{\text{T2I}}$ parameterized by $\theta$ and the text prompt $c$ (i.e., $x_{0}=\Phi_{\text{T2I}}(c;\theta)$). Random Gaussian noise $\epsilon$ is then added to the images based on the pre-defined timestep $t$ and the timestep scheduling coefficient $\bar{\alpha}_{t}$. 111 The $x_{0}$ will be replaced to latent for the latent diffusion models [41].
@@
[…]
@@
where sg denotes the stop-gradient operation. Initially, we quantify concept sensitivity using the RMS norm of the concept loss gradient ($\left\lVert\nabla_{\theta}\mathcal{L}_{\text{Concept}}\right\rVert$) to compare gradients across layers with different dimensions. However, gradients exhibit significant positional bias across layers (Supplementary Section B.1). To address this, we scale the gradient magnitude for each layer by computing the ratio between the gradients of the concept loss and the original diffusion loss ($\left\lVert\nabla_{\theta}\mathcal{L}_{\text{Diffusion}}\right\rVert$), referred to as concept sensitivity.
$\mathcal{L}_{\text{Diffusion}}:=\left\lVert\epsilon_{\theta}(x_{t},c,t)-\epsilon\right\rVert^{2}_{2}$ (6)
$\text{Concept Sensitivity}(\theta):=\mathbb{E}_{x_{0},\epsilon,c_{\text{Aug}}}\left[\frac{\left\lVert\nabla_{\theta}\mathcal{L}_{\text{Concept}}\right\rVert}{\left\lVert\nabla_{\theta}\mathcal{L}_{\text{Diffusion}}\right\rVert}\right]$ (7)
As shown in Eq. 7, we compute the average gradient ratio over generated images ($x_{0}$), added noise ($\epsilon$), and augmented prompts ($c_{\text{Aug}}$). Concept sensitivity varies based on how partial derivatives with respect to each parameter are grouped, such as block-wise or layer-wise. We adopt a projection-wise approach (details in Supplementary Section G.2) to identify weights sensitive to the desired concept, enabling its selective learning by modifying the target of the diffusion loss.
@@ 3.3 Concept-Aware LoRA (CA-LoRA)
Based on the concept sensitivity, we introduce Concept-Aware LoRA (CA-LoRA), a novel parameter-efficient fine-tuning method that selectively updates only concept-aware weights while preserving other weights. CA-LoRA enables the model to learn the desired concept while retaining the pretrained knowledge of T2I models for other concepts.
We begin with Low-Rank Adaptation (LoRA) [21], a widely used fine-tuning method for large-scale pretrained T2I models. For a pretrained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA constrains updates using low-rank decomposition, $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$ and the rank $r\ll\text{min}(d,k)$. In the literature on fine-tuning T2I models, LoRA is commonly applied to all projection layers (Q, K, V, OUT) in each attention block of the pretrained T2I model [41, 21] (left in Fig. 4). LoRA and its extensions [58, 53, 60, 40] aim to improve parameter efficiency for better adaptation to downstream tasks (e.g., by refining low-rank matrix decomposition). However, selective learning approaches for a specific concept from given datasets have not been extensively explored.
CA-LoRA selectively fine-tunes the top-$k$of pretrained weights based on concept sensitivity (Eq. 7), enabling it to learn only the desired concept. Our experiments examine CA-LoRA in terms of viewpoint and style sensitivity. The selected proportions are user-adjustable, allowing control over the extent of fine-tuning, as shown in Fig. 4 (right). In the following sections, we refer to Style CA-LoRA and Viewpoint CA-LoRA as CA-LoRA fine-tuning based on style and viewpoint sensitivity, respectively.
@@ 3.4 Label Generator and Dataset Generation
We train an additional lightweight label generator following DatasetDM [51] to produce a segmentation label corresponding to the image. First, given the real image $x$ from the training set, we add noise to the real image $x$ to synthesize noisy image $x_{t}$. Then, we extract multi-scale generative features (feature maps ${\mathcal{F}}$, cross-attention maps ${\mathcal{A}}$) from the T2I generator $\epsilon_{\theta}(x_{t},c,t)$, which contain lots of semantic information [59, 52, 24]. Last, the Mask2Former-shaped [6] label generator receives the generative features ${\mathcal{F}},{\mathcal{A}}$ and predicts the label maps as illustrated in Stage 3 of Fig. 2. In summary, the entire process can be encapsulated by the text-to-(image-and-label) generator, denoted as $\Phi_{\text{T2(I,L)}}$.
@@
However, a significant domain gap exists between the training and generated images, as shown in Fig. 1. This difference results in generative features at inference exhibiting significantly different statistics from those at training, which in turn degrades the performance of the label generator. Unlike DatasetDM [51], we train the label generator using the fine-tuned T2I model with CA-LoRA. Since fine-tuning the T2I model on training images significantly reduces the domain gap, the label generator receives generative features at inference time that are statistically similar to those at training time, leading to a substantial improvement in image-label alignment.
Generating Diverse Dataset
@@
Lastly, we introduce a method to generate diverse image-label pairs by modifying text prompts. Generating adverse weather conditions (e.g., foggy, night-time) is crucial for improving domain generalization in urban-scene segmentation, as described in DGInStyle [23]. To achieve this, we append weather conditions to the default prompt, e.g., “photorealistic first-person urban street view” to “… in foggy weather”, as illustrated in Stage 4 of Fig. 2. However, as shown in Fig. 1, fine-tuning T2I models with the original LoRA often leads to overfitting to undesired concepts from the source dataset (e.g., clear-day style from Cityscapes [9]) and compromises controllability over textual augmentation. In contrast, our CA-LoRA fine-tunes only the viewpoint-sensitive weights, enabling selective learning of viewpoint concepts while preserving the T2I model’s text adherence in style augmentation. This capability plays a critical role in generating diverse image-label pairs. Additionally, we vary class names in the prompt (e.g., “… with car, person, etc.”). In summary, the generation prompts are formulated as $c_{\text{Gen}}=$ “Photorealistic first-person urban street view with [class names] in [weather condition].”. Now, the generated diverse dataset is utilized together with the real dataset to train segmentation models.
Method Training Dataset Total Iterations Fraction of the Cityscapes Dataset
Real Generated 0.3% 1% 3% 10% 100%
Baseline 90K 41.83 49.15 59.07 69.02 79.40
For a fair comparison, we fine-tune the baseline for additional 30K iterations using real or generated datasets.
Baseline (FT) 120K 42.00 (+0.17) 49.18 (+0.03) 59.06 (-0.01) 68.68 (-0.34) 80.05 (+0.65)
DatasetDM 120K 42.82 (+0.99) 49.71 (+0.56) 60.31 (+1.24) 69.04 (+0.02) 80.45 (+1.05)
Ours 120K 44.13 (+2.30) 51.90 (+2.75) 61.29 (+2.22) 70.29 (+1.27) 80.74 (+1.34)
Table 1: In-domain segmentation performance across various fractions of the Cityscapes dataset (mIoU). In the first row, we trained Mask2Former on various fractions of the Cityscapes dataset (Baseline). Then, we fine-tuned the baseline on DatasetDM and our generated datasets with 30K iterations and evaluated the performance of the fine-tuned segmentation models. Additionally, we include an additional fine-tuned baseline denoted as Baseline (FT), solely fine-tuned on the same real dataset for a fair comparison in terms of the total iterations.
DG Method Generated Dataset ACDC DZ BDD MV Average
ColorAug* 52.38 23.00 53.33 60.06 47.19
ColorAug* DGInStyle 55.19 26.83 55.18 59.95 49.29 (+2.10)
ColorAug 53.12 25.69 53.00 59.81 47.91
@@
ColorAug InstructPix2Pix 56.02 26.92 54.03 60.44 49.35 (+1.44)
ColorAug Ours 56.07 29.75 54.35 61.40 50.39 (+2.49)
DAFormer* 55.15 28.28 54.19 61.67 49.82
DAFormer* DGInStyle 57.74 28.55 56.26 62.67 51.31 (+1.48)
DAFormer 53.98 27.82 54.29 62.69 49.70
@@
DAFormer InstructPix2Pix 55.13 26.93 54.61 62.36 49.76 (+0.06)
DAFormer DATUM† 54.06 27.10 54.74 62.40 49.58 (-0.12)
DAFormer Ours 55.83 31.68 54.68 63.09 51.32 (+1.63)
HRDA* 59.70 31.07 58.49 68.32 54.40
HRDA* DGInStyle 61.00 32.60 58.84 67.99 55.11 (+0.71)
HRDA 58.48 29.46 56.12 64.27 52.08
@@
HRDA InstructPix2Pix 58.50 29.56 56.10 64.10 52.07 (-0.01)
HRDA DATUM† 58.11 30.18 56.94 64.29 52.38 (+0.30)
HRDA Ours 58.93 34.41 56.56 64.54 53.61 (+1.53)
Table 2: Comparison of generated datasets for domain generalization (DG) in urban-scene segmentation (Cityscapes [9] $\rightarrow$ ACDC [45], Dark Zurich [44], BDD100K [57], Mapillary Vistas [33]) (mIoU). The experiments are conducted upon various DG methods (ColorAug [54], DAFormer [19], and HRDA [20]). * The gray color indicates the reported score from the DGInStyle authors. [^3] Due to the absence of the source code with the generated datasets in Cityscapes as a source dataset, we cannot reproduce the reported results. However, we will update the table when the code is available.†DATUM additionally leverages a single target domain image for each weather condition from the ACDC dataset.

### v2

@@ Abstract
Work done during an internship at Qualcomm AI Research.Corresponding author. Work done while at Qualcomm AI Research.Qualcomm AI Research is an initiative of Qualcomm Technologies, Inc.
@@ 1 Introduction
The amount of labeled data is crucial for achieving high performance in semantic segmentation. However, collecting diverse image samples, especially in rare or complex scenarios, and providing pixel-level annotations is both labor-intensive and time-consuming. To address this data scarcity, recent studies [60, 1, 51, 52] have leveraged text-to-image (T2I) generation models [41, 43, 37] to synthesize training data for semantic segmentation. These studies utilize the rich generative capabilities of T2I models to produce promising image-label pairs. In addition, the text-driven controllability of T2I models enables targeted generation of underrepresented distributions. This has proven particularly effective for augmenting labeled datasets, alleviating data scarcity, and generating specific distributions such as balancing class [47] or weather conditions [22].
There are two primary challenges in generating segmentation datasets using image generation models: (1) generating informative datasets beyond the original training data, and (2) achieving alignment with the target domain, as illustrated in Fig. 1. However, prior approaches have often overlooked key aspects of these challenges. Early works on segmentation dataset generation that do not utilize pretrained T2I models [60, 29, 1, 35] typically produce domain-aligned samples, as the generative models are trained solely on the target dataset. Yet, due to the lack of external knowledge, they are limited in their ability to generate informative samples beyond the training distribution.
In contrast, recent studies [51, 52, 57, 22] leverage pretrained T2I models such as Stable Diffusion, which are trained on large-scale datasets like LAION-5B [46], enabling them to incorporate external information. However, these methods often use the pretrained models without task-specific fine-tuning for segmentation, which leads to poor alignment with the target domain.
@@
To address this, we apply existing fine-tuning methods [20] to align the image domain of pretrained generative models with the target domain. However, as illustrated in Figure 1, even LoRA-based fine-tuning tends to overfit and memorize the training data, limiting the model’s ability to generate diverse and informative samples. This overfitting arises because the model learns all concepts present in the training data, such as viewpoint, style, object shape, and layout, regardless of whether they are necessary for domain alignment. To overcome this limitation, we argue that fine-tuning should selectively focus on only the essential concepts (e.g., viewpoint or style) required for domain alignment, while preserving the pretrained T2I model’s capacity to generate informative samples. Moreover, since the relevant concepts for alignment may vary across different problem settings, we propose a flexible fine-tuning method that adapts to the necessary concepts for generating urban-scene segmentation datasets. In in-domain segmentation settings, where the source and target domains are aligned (e.g., clear-day conditions), the T2I model primarily needs to learn the clear-day style of the training data. In contrast, in domain generalization (DG) settings, where the target domain is unseen (e.g., rainy-day conditions), focusing on the driving viewpoint is often more effective, as style alignment is no longer feasible. Our method supports both settings by selectively tuning the relevant concept, enabling the generation of datasets that are both informative and well aligned with the task requirements.
We propose Concept-Aware LoRA (CA-LoRA), a fine-tuning method that automatically identifies and updates only the weights of T2I models associated with the desired concepts, while keeping the remaining weights frozen to preserve the pretrained knowledge. This allows the model to learn only the specified concepts, enabling it to generate images that are both well-aligned with the target domain and more diverse and informative. For example, when the desired concept is the viewpoint, our method fine-tunes the model to learn only the viewpoint from the given dataset, allowing it to generate images with varying styles, object shapes, and layouts—capabilities that standard LoRA fine-tuned models lack, as they tend to overfit to the training style and fail to generalize beyond it, even with varied text prompts. Consequently, our fine-tuned T2I models can generate specific styles based on user input, making them particularly effective in DG settings [6, 19, 18].
@@
Our analysis shows that the proposed CA-LoRA can exclusively learn the specified concepts, and we demonstrate the effectiveness of the generated datasets for urban-scene segmentation by comparing them with state-of-the-art dataset generation methods [51, 22, 4, 3] in both in-domain and domain generalization settings.
Our contributions are threefold:
@@
    We propose Concept-Aware LoRA, a novel fine-tuning method that selectively identifies and updates only the weights related to the necessary concepts, reducing overfitting and preserving pretrained knowledge.
@@
    Our method achieves state-of-the-art performance across various settings, including a +2.30% mIoU improvement in few-shot and +1.34% in fully supervised settings on Cityscapes, as well as an average gain of +1.53% mIoU on domain generalization benchmarks.
Figure 2: Overview of our framework for generating an urban-scene segmentation dataset by learning the Cityscapes viewpoint. The process consists of four stages: (1) identifying sensitive weights for a specific concept, (2) selectively fine-tuning them with LoRA, (3) training a label generator using features from T2I model, and (4) generating diverse image-label pairs with augmented prompts.
@@ 2.2 Segmentation Dataset Generation
Recent studies have focused on leveraging the pretrained knowledge of T2I models [51, 52, 34, 3, 13, 23], but often overlook the alignment of the generated images with the target domain (e.g., style, viewpoint). This paper explores the impact of fine-tuning T2I models for segmentation dataset generation, focusing on ensuring better image alignment.
Figure 3: Overview of measuring concept awareness. (a) We design the concept loss ($\mathcal{L}_{\text{Concept}}$) with the concept-augmented captions ($c_{\text{Aug}}$), and the original diffusion loss ($\mathcal{L}_{\text{Diff}}$) with the added noise $\epsilon$. The concept-augmented captions can be changed according to the desired concept (e.g., style, viewpoint). (b) While each concept gradient represents the reaction of the concept, it has to be normalized with the original diffusion gradient to assess the increased ratio of each layer.
@@ 3.1 Overall Framework
Figure 2 shows our dataset generation framework. First, we identify weights sensitive to the desired concept to be learned from the source dataset (Stage 1). Next, Concept-Aware LoRA selectively fine-tunes the top-kk% sensitive weights of the T2I model (Stage 2). Then, a label generator is trained on a labeled dataset using rich generative features from the T2I model (Stage 3). Finally, diverse image-label pairs are generated with augmented prompts according to the problem settings (Stage 4).
@@ 3.2 Identify Sensitive Weights to Desired Concept
To identify the sensitive weights for a specific concept (e.g., style, viewpoint), we design an objective function, Concept loss ($\mathcal{L}_{\text{Concept}}$), which can be flexibly changed according to the desired concept. As shown in Figure 3 (a), Concept loss is applied to the noisy image to enforce modifying the concept. For example, when the Concept loss forces the T2I model to modify style (e.g., photorealistic $\rightarrow$ sketch), its gradient can be used to identify style-aware weights.
First, for the Concept loss input, we generate a few images $x_{0}$ using the T2I model $\Phi_{\text{T2I}}$ parameterized by $\theta$ and the text prompt $c$ (i.e., $x_{0}=\Phi_{\text{T2I}}(c;\theta)$). Random Gaussian noise $\epsilon$ is then added to the images based on the pre-defined timestep $t$ and the timestep scheduling coefficient $\bar{\alpha}_{t}$.
@@
$c=\text{``Photorealistic first-person urban street view''},$ (2)
@@
where sg denotes the stop-gradient operation. Initially, we quantify concept awareness using the RMS norm of the concept loss gradient ($\left\lVert\nabla_{\theta}\mathcal{L}_{\text{Concept}}\right\rVert$) to compare gradients across layers with different dimensions. However, the gradients show a significant positional bias across layers (see Appendix B for details). To address this, we scale the gradient magnitude for each layer by computing the ratio between the gradients of the concept loss and the original diffusion loss ($\left\lVert\nabla_{\theta}\mathcal{L}_{\text{Diff}}\right\rVert$), referred to as concept awareness.
$\mathcal{L}_{\text{Diff}}:=\left\lVert\epsilon_{\theta}(x_{t},c,t)-\epsilon\right\rVert^{2}_{2}$ (6)
$\text{Concept-Awareness}(\theta):=\mathbb{E}_{x_{0},\epsilon,c_{\text{Aug}}}\left[\frac{\left\lVert\nabla_{\theta}\mathcal{L}_{\text{Concept}}\right\rVert}{\left\lVert\nabla_{\theta}\mathcal{L}_{\text{Diff}}\right\rVert}\right]$ (7)
We average the ratio over images ($x_{0}$), noise ($\epsilon$), and augmented prompts ($c_{\text{Aug}}$) to ensure statistical robustness.
Using the measured concept awareness values computed for each weight, we group and average them by attention projection layers (Q, K, V, and OUT) to quantify the concept awareness of each projection layer in the T2I model.
@@ 3.3 Concept-Aware LoRA (CA-LoRA)
After identifying concept-sensitive weights, we then apply LoRA updates only to these selected layers. Concept-Aware LoRA (CA-LoRA), our parameter-efficient fine-tuning method, selectively updates only the concept-relevant weights while keeping the remaining parameters frozen. This selective adaptation allows the model to learn the desired concept, yet still retain the pretrained T2I model’s knowledge across other, unrelated concepts.
@@
We begin with Low-Rank Adaptation (LoRA) [20], a widely used fine-tuning method for large-scale pretrained T2I models. For a pretrained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA constrains updates using low-rank decomposition, $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$ and the rank $r\ll\text{min}(d,k)$. In the literature on fine-tuning T2I models, LoRA is commonly applied to all projection layers (Q, K, V, OUT) in each attention block of the pretrained T2I model [41, 20] (left in Figure 4). LoRA and its extensions [59, 53, 61, 40] aim to improve parameter efficiency for better adaptation to downstream tasks (e.g., by refining low-rank matrix decomposition). However, selective learning approaches for a specific concept from given datasets have not been extensively explored.
CA-LoRA selectively fine-tunes the top-kk% of projection layers based on concept awareness (Eq. 7), enabling it to learn only the desired concept. Our experiments examine CA-LoRA in terms of viewpoint and style awareness. The selected proportions are user-adjustable, allowing control over the extent of fine-tuning, as shown in Figure 4 (right). In the following sections, we refer to Style CA-LoRA and Viewpoint CA-LoRA as CA-LoRA fine-tuning based on style and viewpoint awareness, respectively.
@@ 3.4 Label Generator and Dataset Generation
We train an additional lightweight label generator following DatasetDM [51] to produce a segmentation label corresponding to the image. First, given the real image $x$ from the training set, we add noise to the real image $x$ to synthesize noisy image $x_{t}$. Then, we extract multi-scale generative features (feature maps $\mathcal{F}$}, cross-attention maps $\mathcal{A}$}) from the T2I generator $\epsilon_{\theta}(x_{t},c,t)$, which contain lots of semantic information [60, 52, 24]. Last, the Mask2Former-shaped [5] label generator receives the generative features ${\mathcal{F}},{\mathcal{A}}$ and predicts the label maps as illustrated in Stage 3 of Figure 2. In summary, the entire process can be encapsulated by the text-to-(image-and-label) generator, denoted as $\Phi_{\text{T2(I,L)}}$.
Reducing Domain Gap in Generative Feature
@@
Importantly, a notable domain gap exists between the training images and the generated images from the pretrained T2I models, as shown in Figure 1. This discrepancy leads to a domain gap in the generative features at inference, whose statistics differ substantially from those at training, thereby degrading the performance of the label generator. Unlike DatasetDM [51], we train the label generator using the fine-tuned T2I model with CA-LoRA. Fine-tuning the T2I model on the training images greatly reduces this domain gap, allowing the label generator to receive generative features at inference that are statistically consistent with those seen during training, which in turn results in a substantial improvement in image-label alignment.
Generating Diverse Image-Label Paired Dataset
@@
Generating adverse weather or light conditions (e.g., foggy, night-time) is crucial for improving domain generalization in urban-scene segmentation, as described in DGInStyle [22]. To achieve this, we append weather conditions as illustrated in Stage 4 of Figure 2. However, fine-tuning T2I models with the original LoRA often leads to overfitting to undesired concepts from the source dataset (e.g., clear-day style) and compromises controllability over textual augmentation. In contrast, our CA-LoRA fine-tunes only the viewpoint-aware layers, enabling selective learning of viewpoint concepts while preserving the T2I model’s text adherence in style augmentation. This capability plays a critical role in generating diverse image-label pairs. Also, we vary class names in the prompt (e.g., “… with car, person, etc.”) to provide additional diversity. In summary, the generation prompts are formulated as $c_{\text{Gen}}=$ “Photorealistic first-person urban street view with [class names] in [weather].”.
Table 1: In-domain segmentation performance across various fractions of the Cityscapes dataset (mIoU). The first row presents the baseline model trained solely on the real dataset. We visualize the performance improvement relative to the baseline alongside each score.
Method Fraction of the Cityscapes Dataset
0.3% 1% 3% 10% 100%
Baseline 41.83 49.15 59.07 69.02 79.40
InstructPix2Pix 41.94 (+0.11) 48.17 (-0.98) 60.43 (+1.36) 66.21 (-2.81) 78.06 (-1.34)
DatasetDM 42.82 (+0.99) 49.71 (+0.56) 60.31 (+1.24) 69.04 (+0.02) 80.45 (+1.05)
LoRA 42.97 (+1.14) 51.80 (+2.65) 60.22 (+1.15) 69.21 (+0.19) 79.75 (+0.35)
AdaLoRA 43.67 (+1.84) 48.21 (-0.94) 60.93 (+1.86) 68.32 (-0.70) 78.62 (-0.78)
CA-LoRA (Ours) 44.13 (+2.30) 51.90 (+2.75) 61.29 (+2.22) 70.29 (+1.27) 80.74 (+1.34)
Table 2: Comparison of generated datasets for domain generalization (DG) of urban-scene segmentation (mIoU). The experiments are conducted using various DG methods [54, 18, 19]. Each first row presents the baseline model trained solely on the real dataset.
DG Method Method ACDC DZ BDD MV Average
Baseline 53.12 25.69 53.00 59.81 47.91
InstructPix2Pix 56.02 26.92 54.03 60.44 49.35 (+1.44)
DATUM 54.47 29.25 53.19 61.22 49.53 (+1.61)
@@
LoRA 54.25 28.42 54.34 61.42 49.61 (+1.70)
AdaLoRA 54.39 28.65 53.78 61.78 49.65 (+1.74)
CA-LoRA (Ours) 56.07 29.75 54.35 61.40 50.39 (+2.49)
Baseline 53.98 27.82 54.29 62.69 49.70
InstructPix2Pix 55.13 26.93 54.61 62.36 49.76 (+0.06)
DATUM 54.06 27.10 54.74 62.40 49.58 (-0.12)
@@
LoRA 54.64 30.22 55.44 63.39 50.92 (+1.22)
AdaLoRA 54.94 30.06 55.41 63.10 50.88 (+1.18)
CA-LoRA (Ours) 55.83 31.68 54.68 63.09 51.32 (+1.63)
Baseline 58.48 29.46 56.12 64.27 52.08
InstructPix2Pix 58.50 29.56 56.10 64.10 52.07 (-0.01)
DATUM 58.11 30.18 56.94 64.29 52.38 (+0.30)
@@
LoRA 57.47 31.09 57.29 64.47 52.58 (+0.50)
AdaLoRA 58.37 32.23 56.40 64.58 52.90 (+0.82)
CA-LoRA (Ours) 58.93 34.41 56.56 64.54 53.61 (+1.53)